1.3g of glycerol, C3H8O3,is mixed with water to form a solution with a volume of 0.25L at 35oC. What is the osmotic pressure (in atmospheres) of the solution.
Answer #1

Alycerol (C3H803) = 1.39m MMCZ HgO, = 3x12.0) + 8 (1.008) + 3(16) ve 0-25 lit - 92.09 gmimo! T= 35€ = 35+273-15 nca HgO3 = 13